# Sequential estimation of disturbed aerodynamic flows from sparse measurements via a reduced latent space

### v1

@@ Abstract
This work presents a fast, scalable, and uncertainty-aware methodology suitable for real-time estimation of key aerodynamic states—including instantaneous vorticity fields and aerodynamic loads—during severe gust encounters, where strong flow separation and vortex–gust interactions dominate the dynamics. The methodology, trained and tested on high-fidelity simulations of two-dimensional wing-gust encounters and corresponding sparse pressure data, combines the strengths of physics-based ensemble filtering with data-driven surrogate modeling in a learned low-order space. Within this reduced space, latent dynamics and sensor observation operators are learned using neural networks, enabling efficient and expressive modeling of complex aerodynamic responses. These data-driven components are then embedded within a low-rank Ensemble Kalman Filter (LREnKF), yielding a computationally efficient scheme that combines data-driven expressivity with physical interpretability in the original high-dimensional space. Because assimilation occurs entirely within the learned reduced space, the updates are fast enough for real-time use in flight applications, ensuring that aerodynamic states can be continuously estimated from streaming pressure data. The filtering process corrects the predicted states predominantly in directions that are both dynamically significant and observable, as determined by the dominant modes of the learned observation operator. This targeted correction mechanism improves filter efficiency and mitigates the risk of over- or under-adjusting state variables in unobservable or weakly constrained subspaces. An analysis of the observability reveals how sensor informativeness evolves over time, particularly during wing-gust interactions. The framework readily accounts for sensor failure, and sensor dropout experiments show that LREnKF adaptively re-weights neighboring sensors to compensate for lost information, preserving estimation quality even under degraded sensing configurations.
@@ 1 Introduction
Aerial disturbances, particularly those arising from various gust encounters, frequently disrupt the nominal operation and controlled maneuvering of small air vehicles. In gust-passage scenarios, the disturbance acts over a limited time window, creating sharp, transient deviations in the flow field that are especially difficult to detect and predict in real time. These disruptions can be mitigated through active control strategies—such as linear feedback control or reinforcement learning (RL)—that enable the system to respond adaptively. However, effective active control requires real-time access to the underlying states of the dynamical system to detect anomalies and take corrective action. In aerodynamic applications, these states are typically not directly measurable. Instead, transient flow variations are reflected in (and partially observable through) the sensors, such as pressure taps mounted on the wings. These measurements serve as indirect observables that can be sequentially assimilated to infer the hidden flow states. Without a good initial guess, it is impossible to infer the inherently high-dimensional flow field from a single measurement taken from a limited number of sensors. Though sequential filters, such as those as the heart of this work, leverage past measurements and dynamical predictions to progressively narrow the search region, each measurement update of the state is still heavily challenged to navigate the high-dimensional space.
@@
To address this challenge, a low-dimensional state space can be constructed to serve as a compressed representation of the full flow field, describing only the most energetic structures and essential physics. Traditional approaches rely on linear projection techniques, such as Proper Orthogonal Decomposition (POD) (Berkooz et al., 1993) and Dynamic Mode Decomposition (DMD) (Schmid, 2010), which approximate the system in terms of dominant linear modes. However, these methods have been shown to be inadequate for strongly nonlinear systems—such as gust-disturbed flows—where the underlying dynamics often evolve on a nonlinear manifold (Fukami and Taira, 2023). In such settings, the modes excluded by linear truncation can significantly affect the system’s evolution (Maulik et al., 2020). To overcome this limitation, nonlinear data compression techniques based on deep learning (DL)—such as standard autoencoders (Fukami and Taira, 2023, Xie et al., 2024, Mousavi and Eldredge, 2025) and their variational counterparts (VAEs) (Fraccaro et al., 2017, He et al., 2025)—have gained traction for their ability to learn compact, task-specific latent representations that better capture the transient dynamics of complex systems.
@@
Once a reduced-order latent space is learned, data assimilation (DA) can be performed efficiently within this space to enable real-time state estimation with low computational overhead using sequential filtering. From the estimated latent state and its uncertainty, the corresponding estimates in the full flow field state space can be reconstructed via the decoder. Typically, filtering frameworks involve two critical components to evolve the probability distribution of states: a forecast operator that propagates states forward in time, and an observation operator that maps states to measurable outputs, enabling state correction based on new observations (Asch et al., 2016, Eldredge and Mousavi, 2025). In a reduced space, these operators can be approximated by DL-based surrogate models, providing a tractable alternative to traditional high-fidelity solvers (Bach et al., 2024). With the abundance of data, neural networks (NNs), in particular, are well-suited for this task as they are differentiable and can either augment or fully replace physics-based models in scenarios where simulations are expensive or unavailable (Jin et al., 2021).
@@
Several recent works have demonstrated the successful use of NNs for learning surrogate models. For instance, Fraccaro et al., (2017) developed the Kalman Variational Autoencoder (Kalman-VAE), which uses a recurrent neural network (RNN) to model latent dynamics from partially observed image sequences. Maulik et al., (2020) employed long short-term memory (LSTM) networks to model the evolution of unresolved POD modes in turbulent flows. In a Kalman filtering setting, Coskun et al., (2017) used three separate LSTM models to jointly learn the latent dynamics and the associated process and measurement noise statistics. Other works embed physical constraints into the learned latent dynamics: Popov and Sandu, (2022) and He et al., (2025) constructed physics-informed surrogate models constrained by PDE structure. The sequence prediction models employed in the foregoing references leverage the full temporal history to forecast future states, often using RNNs with memory kernels (Tang et al., 2020, Cheng et al., 2022).
Despite their effectiveness, memory-based models violate the foundational Markov assumption in filtering, which posits that system evolution depends only on the current states (Maybeck, 1982). To preserve this assumption, Neural Ordinary Differential Equations (Neural ODEs) have emerged as a promising alternative. Neural ODEs define continuous-time latent dynamics via:
$\frac{\mathrm{d}x}{\mathrm{d}t}=f_{\theta}(x)$ (1)
where $f_{\theta}(x)$ is a NN parameterized by $\theta$. Unlike RNNs, Neural ODEs inherently model Markovian transitions, making them particularly compatible with sequential filtering architectures.
@@
Building on this foundation, several studies have integrated DL techniques into traditional filtering frameworks for improved state estimation. Jouaber et al., (2021) introduced a neural Kalman filter in which an LSTM dynamically learns the process noise covariance, with the full Kalman update embedded in the computational graph. Song et al., (2022) proposed KFFLSTM, a hybrid architecture that jointly learns dynamics and noise models using LSTMs, though it departs from the Markovian assumption. Revach et al., (2022) developed KalmanNet, which learns the Kalman gain in extended Kalman filtering with known models but unknown noise statistics, albeit without explicit uncertainty quantification. In large-scale geophysical applications, Wang and Shen, (2024) used NNs to enhance ensemble adjustment Kalman filters for strongly coupled ocean–atmosphere systems, and Howard et al., (2024) developed a CNN-based surrogate to accelerate the Ensemble Kalman Filter (EnKF) analysis step in high-resolution observational settings. Besides model-based filtering, fully model-free Bayesian estimators have also been introduced: Ghosh et al., 2024b proposed DANSE, an RNN-based Bayesian filter that infers posterior distributions from noisy observations, and Ghosh et al., 2024a extended it to a semi-supervised version, SemiDANSE, incorporating state supervision as a regularization term. A recent line of work (Özalp et al., 2025) has addressed DA by coupling convolutional autoencoders with recurrent latent models and ensemble Kalman filtering, resulting in the DA-CAE-ESN framework for chaotic flows such as Kuramoto–Sivashinsky and Kolmogorov turbulence. Beyond filtering, variational DA approaches like 3D-Var and 4D-Var remain widely used for deterministic state estimation via optimization of forecast-observation mismatches (Le Dimet and Talagrand, 1986, Kalnay, 2003, de Rosnay et al., 2022, Bach et al., 2024). However, these methods only yield point estimates and lack real-time adaptability, making them unsuitable for control-driven aerodynamic applications requiring fast and probabilistic updates.
This work introduces a data assimilation algorithm for unsteady fluid flows in a reduced-dimensional latent space. The framework we present provides real-time flow state estimation in aerodynamic environments with strong transient disturbances, observed only with sparse and noisy surface pressure measurements. While our prior work (Mousavi and Eldredge, 2025) explored the inference of the latent state and its uncertainty from sparse, noisy measurements, it considered this inference only as a single instantaneous task and did not leverage the evolving state estimate from past observations. The present study fills this gap by embedding DA within a learned latent space, enabling sequential state estimation conditioned on the full history of observations. The framework employs a nonlinear physics-augmented autoencoder to extract a compact latent representation of the high-dimensional vorticity field. The latent state is updated in time using sequential filtering, in which both the forecast and observation operators are learned via neural networks and incorporated into the low-rank ensemble Kalman filter (LREnKF) introduced by Le Provost et al., (2022). Unlike the DA-CAE-ESN framework of Özalp et al., (2025), which was demonstrated on canonical chaotic flows through augmented reservoir state and reported limited estimation accuracy even with sixteen sensors, our method is specifically designed for aerodynamic gust encounters and achieves robust reconstructions from substantially fewer measurements. Through the LREnKF formulation used in this study, updates are restricted to measurement-informative latent directions, enabling efficient use of very sparse sensing. Thus, this integration of the low-dimensional latent space with a structured sequential estimator ensures both computational efficiency and physical interpretability in the uncertainty-aware estimation of the evolving aerodynamic state under gust-induced disturbances.
@@ 2 Problem description and methodology
In this study, we present a unified framework that integrates data-driven modeling with sequential estimation to reconstruct unsteady flow fields and aerodynamic loads from sparse, noisy surface pressure measurements. The central aim is to enable fast, robust, and uncertainty-aware inference under strongly disturbed conditions, where traditional simulation or filtering approaches become computationally prohibitive or inaccurate. Figure 1 illustrates the overall concept. At the heart of the framework is a compact latent space, $\boldsymbol{\xi}$, learned through nonlinear compression techniques based on DL. This reduced representation captures the essential flow physics while filtering out redundancy, thereby providing a tractable foundation for data assimilation. Surrogate dynamical and observation models—required by DA task—are then trained directly in this latent space. During inference, new surface pressure data are assimilated with the LREnKF, which updates the latent state distribution efficiently by exploiting measurement-informative subspaces. Finally, the updated latent samples are decoded back into the physical space, yielding flow-field reconstructions and load estimates along with their associated uncertainties.
@@ 2.1 Problem statement
We need data for DL tasks. To generate training and testing data, we simulate an unsteady, incompressible two-dimensional flow over a NACA 0012 airfoil at angles of attack (AoA) $\alpha\in\{20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ}\}$. The undisturbed flow is characterized by a chord-based Reynolds number of $\mathrm{Re}=U_{\infty}c/\nu=100$, where $U_{\infty}$ is the freestream velocity, $c$ is the chord length, and $\nu$ is the kinematic viscosity of the fluid. To model disturbed conditions, we introduce localized gusts upstream of the airfoil at randomly selected instants during the natural vortex shedding cycle of the airfoil. These gusts are implemented as a body force $\boldsymbol{F}$—Gaussian distributed in space and time, and applied to the flow field—defined as:
@@
The direct numerical simulation of the incompressible Navier–Stokes equations is performed in vorticity–streamfunction form using the Lattice Green’s Function/Immersed Layers method developed by Eldredge, (2022). This approach enables high-resolution flow computation within a relatively compact domain, made possible by the efficient treatment of boundary conditions via the lattice Green’s function. The simulations are conducted over a Cartesian domain spanning $(-2c,4c)\times(-2c,2c)$ in the $x$ and $y$ directions, respectively, with uniform grid spacing $\Delta x/c=0.02$. During the simulation, surface pressure $p_{s}$ relative to the ambient pressure $p_{\infty}$ is recorded at 11 predefined sensor locations (see figure 2). The pressure coefficient is computed using the standard definition:
$C_{p}=\frac{2(p_{s}-p_{\infty})}{\rho U_{\infty}^{2}}.$ (3)
For training purposes, the vorticity data are extracted from a tighter region within the computational domain: $(-1.5c,3.3c)\times(-1.2c,1.2c)$ in the $x$ and $y$ directions, respectively. This subdomain is selected to maintain a balance between computational efficiency and physical fidelity, ensuring the full evolution of the gust disturbance and its interaction with the airfoil and wake are accurately captured. At each AoA, one undisturbed (baseline) case and 60 randomized gust cases are simulated, resulting in a dataset comprising 5 undisturbed cases and 300 disturbed (each with a single gust passage) cases in total. Each case is computed over 10 nondimensional convective time units, defined as $t=t^{\prime}U_{\infty}/c$, where $t^{\prime}$ is the dimensional time. This time horizon ensures that at least two full vortex shedding cycles are captured. Flow fields are sampled at a uniform time interval of $\Delta t=0.02$, yielding 500 snapshots per case. Altogether, the dataset consists of 152500 flow snapshots (vorticity field, surface pressure at the designated sensor locations, and lift force) spanning a wide range of aerodynamic conditions. The flow system states are defined as the vorticity field and the corresponding lift coefficient, which are to be estimated by assimilating synthetic surface pressure measurements.

### v2

@@ Abstract
This work presents a fast and uncertainty-aware sequential data assimilation framework suitable for estimation of key aerodynamic states (e.g., instantaneous vorticity fields and aerodynamic loads) during severe gust encounters, where vortex–gust interactions strongly affect the flow dynamics. The framework comprises an ensemble Kalman filter (EnKF), designed to detect and reconstruct nearly-impulsive flow disturbances with a wide range of strengths and orientations and introduced at arbitrary times. The forecast and measurement update (analysis) stages of the EnKF are each composed of learned operators in a low-dimensional latent space, obtained via a physics-augmented autoencoder. The forecast operator propagates the undisturbed baseline dynamics, but cannot predict random gust-induced deviations from these dynamics. Thus, the analysis stage frequently assimilates new surface pressure measurements to listen for disturbance signals and initiate deviation from the nominal trajectory. The methodology is trained and tested on flowfield snapshots from high-fidelity simulations of two-dimensional airfoil-gust encounters and corresponding sparse surface pressure data. Because assimilation occurs entirely within the reduced-order latent space, the updates are computationally efficient and ensure that aerodynamic states can be continuously estimated from streaming pressure data. The estimated instantaneous latent state remains physically interpretable via decoding to the original high-dimensional flow. Eigenvalue decomposition of the state and observation Gramians in the measurement update stage reveals the dominant correction directions necessary to capture the flow disturbance and quantifies how sensors inform the state corrections throughout gust interaction. The framework readily accounts for sensor failure; sensor dropout experiments show that the EnKF adaptively re-weights neighboring sensors to compensate for lost information, preserving estimation quality even under degraded sensing configurations.
@@ 1 Introduction
Aerial disturbances, particularly those arising from various gust encounters, frequently disrupt the nominal operation and controlled maneuvering of small air vehicles. In gust-passage scenarios, the disturbance acts over a limited time window, creating sharp, transient deviations in the flow field that are especially difficult to detect and predict in real time. These disruptions can be mitigated through active control strategies—such as linear feedback control or reinforcement learning (RL)—that enable the system to respond adaptively [28]. However, effective active control requires real-time access to the underlying states of the dynamical system to detect anomalies and take corrective action. In aerodynamic applications, these states are typically not directly measurable. Instead, transient flow variations are reflected in (and partially observable through) the surface sensors, such as pressure taps mounted on the wings. These measurements serve as indirect observables that can be sequentially assimilated to infer the hidden flow states. In such scenarios, without a good initial guess, it is impossible to infer the inherently high-dimensional flow field from a single measurement taken from a limited number of sensors. Though sequential filters, such as those as the heart of this work, leverage past measurements and dynamical predictions to progressively narrow the search region, each measurement update of the state is still heavily challenged to navigate the high-dimensional space.
@@
To address this challenge, a low-dimensional state space can be constructed to serve as a compressed representation of the full flow field, describing only the most energetic structures and essential physics. Traditional approaches rely on linear projection techniques, such as Proper Orthogonal Decomposition (POD) [4, 23] and Dynamic Mode Decomposition (DMD) [39], which approximate the system in terms of dominant linear modes. However, these methods have been shown to be inadequate for classes of strongly nonlinear systems—such as gust-disturbed flows—where the underlying dynamics often evolve on a nonlinear manifold [15]. In such settings, the modes excluded by linear truncation can significantly affect the system’s evolution [29]. For gust–disturbed flow over an airfoil at low Reynolds number ($Re=100$), Fukami and Taira [15] demonstrated that the POD coordinates in the learned latent space cause dynamically distinct vorticity regimes to collapse onto overlapping trajectories. In contrast, the same study demonstrated that nonlinear autoencoders are capable of constructing latent representations that preserve the dynamical distinctions between flow states, enabling accurate low-dimensional compression and reconstruction of transient gust–airfoil interactions. Motivated by this improved expressiveness, nonlinear data compression techniques based on deep learning (DL), such as standard autoencoders [15, 44, 32] and their variational counterparts (VAEs) [13, 18], have gained traction for their ability to learn compact, task-specific latent representations that more faithfully capture the transient dynamics of complex flow systems.
@@
Once a reduced-order latent space is learned, data assimilation (DA) can be performed efficiently within this space to enable fast posterior state estimation with low computational overhead using sequential filtering. From the estimated latent state and its uncertainty, the corresponding estimates in the full flow field state space can be reconstructed via the decoder. Typically, filtering frameworks involve two critical components to evolve the probability distribution of states: a forecast operator that propagates states forward in time, and an observation operator that maps states to measurable outputs, enabling state correction based on new observations [1, 10]. In a reduced space, these operators can be approximated by DL-based surrogate models, providing a tractable alternative to traditional high-fidelity solvers [2]. With the abundance of data, neural networks (NNs), in particular, are well-suited for this task as they are designed to be differentiable, enabling gradient-based optimization, and they can either augment or fully replace physics-based models in scenarios where simulations are expensive or unavailable [20].
@@
Several recent works have demonstrated the successful use of NNs for learning surrogate models. For instance, Fraccaro et al. [13] developed the Kalman Variational Autoencoder (Kalman-VAE), which uses a recurrent neural network (RNN) to model latent dynamics from partially observed image sequences. Maulik et al. [29] employed long short-term memory (LSTM) networks to model the evolution of unresolved POD modes in turbulent flows. In a Kalman filtering setting, Coskun et al. [6] used three separate LSTM models to jointly learn the latent dynamics and the associated process and measurement noise statistics. Other works embed physical constraints into the learned latent dynamics: Popov and Sandu [36] and He et al. [18] constructed physics-informed surrogate models constrained by PDE structure. The sequence prediction models employed in the aforementioned references leverage the full temporal history to forecast future states, often using RNNs with memory kernels [42, 5].
@@
Other studies have integrated DL techniques into traditional filtering frameworks for improved state estimation. Jouaber et al. [21] introduced a neural Kalman filter in which an LSTM dynamically learns the process noise covariance, with the full Kalman update embedded in the computational graph. Song et al. [41] proposed KFFLSTM, a hybrid architecture that jointly learns dynamics and noise models using LSTMs. Revach et al. [37] developed KalmanNet, which learns the Kalman gain in extended Kalman filtering with known models but unknown noise statistics, albeit without explicit uncertainty quantification. In large-scale geophysical applications, Wang and Shen [43] used NNs to enhance ensemble adjustment Kalman filters for strongly coupled ocean–atmosphere systems, and Howard et al. [19] developed a CNN-based surrogate to accelerate the ensemble Kalman filter (EnKF) analysis step in high-resolution observational settings. Besides model-based filtering, fully model-free Bayesian estimators have also been introduced: Ghosh et al. [17] proposed DANSE, an RNN-based Bayesian filter that infers posterior distributions from noisy observations, and they further extended it to a semi-supervised version, SemiDANSE [16], incorporating state supervision as a regularization term. A recent line of work by Özalp et al. [33] and Ozan et al. [35, 34] has advanced data assimilation and control of chaotic flows by integrating nonlinear autoencoders, latent dynamical models, filtering, and reinforcement learning, resulting in unified frameworks for forecasting and control in partially observed chaotic systems. Beyond filtering, variational DA approaches like 3D-Var and 4D-Var remain widely used for deterministic state estimation via optimization of forecast-observation mismatches [24, 22, 9, 2]. However, these variational methods yield only point estimates and lack online adaptability, making them unsuitable for control-driven aerodynamic applications requiring fast and probabilistic updates. While these approaches provide powerful filtering frameworks, they have not been developed with the specific challenges of aerodynamic gust detection, including high-dimensional flow states, sparse pressure sensing, and disturbance onset occurring at unknown times.
The current work introduces a data assimilation framework for unsteady aerodynamic flows, designed to detect and reconstruct sudden gust encounters from sparse surface pressure measurements. Each encounter originates from the same undisturbed baseline flow state, and because gust disturbances are introduced at arbitrary times, a learned forecast model alone cannot anticipate the transition from the nominal baseline trajectory to a gust-driven response without a priori knowledge of the gust parameters. In the absence of measurement updates, the forecast would simply continue to propagate the undisturbed dynamics. The central contribution of this study is the systematic use of the Kalman filtering framework—specifically its analysis step—to leverage surface pressure sensors as transient disturbance detectors that identify and correct for gust encounters of varying strength and orientation occurring at arbitrary times. Through frequent measurement updates, the filter continuously assimilates sensor information, enabling rapid detection of deviations from the baseline flow and reconstruction of the full aerodynamic response. This sequential estimation mechanism is applied to a compact latent representation of the high-dimensional vorticity field extracted by a nonlinear physics-augmented convolutional autoencoder. Both the forecast and observation operators are learned in the latent space and incorporated into the ensemble Kalman filter (EnKF), enabling computationally efficient updates while preserving a physically interpretable uncertainty structure. Further eigendecomposition of the state and measurement Gramians enables quantification of the instantaneous informativeness of individual sensors as gust vortices convect over the airfoil and interact with vortices from the wing and in the wake. The resulting methodology provides scalable, uncertainty-aware reconstruction of gust-driven aerodynamic dynamics from sparse measurements.
@@ 2 Problem description and methodology
In this study, we present a unified framework that integrates data-driven modeling with sequential estimation to reconstruct unsteady flow fields and lift from sparse, noisy surface pressure measurements. The central aim is to enable fast, robust, and uncertainty-aware inference under strongly disturbed conditions, where traditional simulation or filtering approaches become computationally prohibitive or inaccurate. Figure 1 illustrates the overall concept. At the heart of the framework is a compact latent space, $\boldsymbol{\xi}$, learned through nonlinear compression techniques based on DL. This reduced representation captures the essential flow physics while filtering out redundancy, thereby providing a tractable foundation for data assimilation. Surrogate dynamical and observation models—required by the DA task—are then trained directly in this latent space. During inference, new surface pressure data are assimilated with the EnKF. Finally, the updated latent samples are decoded back into the physical space, yielding flow-field reconstructions and load estimates along with their associated uncertainties. In this section, we describe the details of the generation of training data and the estimation framework itself.
@@ 2.1 Problem statement
To generate training and testing data for DL tasks, we simulate an unsteady, incompressible two-dimensional flow over a NACA 0012 airfoil at fixed angles of attack (AoA), $\alpha\in\{20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ}\}$. The undisturbed flow is characterized by a chord-based Reynolds number of $\mathrm{Re}=U_{\infty}c/\nu=100$, where $U_{\infty}$ is the freestream velocity, $c$ is the chord length, and $\nu$ is the kinematic viscosity of the fluid. To model disturbed conditions, we introduce localized gusts upstream of the airfoil at randomly selected instants during the natural vortex shedding cycle of the airfoil. These gusts are implemented as a body force $\boldsymbol{F}$—Gaussian distributed in space and time, and applied to the flow field—defined as:
@@
The direct numerical simulation of the incompressible Navier–Stokes equations is performed in vorticity–streamfunction form using the Lattice Green’s Function/Immersed Layers method developed by Eldredge [11]. This approach enables high-resolution flow computation within a relatively compact domain, made possible by the efficient treatment of boundary conditions via the lattice Green’s function. The simulations are conducted over a Cartesian domain spanning $(-2c,4c)\times(-2c,2c)$ in the $x$ and $y$ directions, respectively, with uniform grid spacing $\Delta x/c=0.02$. During the simulation, surface pressure $p_{s}$ relative to the ambient pressure $p_{\infty}$ is recorded at 11 predefined uniformly-distributed sensor locations (see figure 2). The pressure coefficient is computed using the standard definition:
$p=\frac{2(p_{s}-p_{\infty})}{\rho U_{\infty}^{2}}.$ (2)
For training purposes, the vorticity data are extracted from a tighter region within the computational domain: $(-1.5c,3.3c)\times(-1.2c,1.2c)$ in the $x$ and $y$ directions, respectively. This subdomain is selected to maintain a balance between computational efficiency and physical fidelity, ensuring the full evolution of the gust disturbance and its interaction with the airfoil and wake are accurately captured. At each AoA, one undisturbed (baseline) case and 60 randomized gust cases are simulated, resulting in a dataset comprising 5 undisturbed cases and 300 disturbed (each with a single gust passage) cases in total. Each case is computed over 10 nondimensional convective time units, defined as $t=t^{\prime}U_{\infty}/c$, where $t^{\prime}$ is the dimensional time. This time horizon ensures that at least two full vortex shedding cycles are captured. Flow fields are sampled at a uniform time interval of $\Delta t=0.02$, yielding 500 snapshots per case. Altogether, the dataset consists of 152 500152\,500 flow snapshots (vorticity field, surface pressure at the designated sensor locations, and lift force) spanning a wide range of aerodynamic conditions. The flow states are defined as the vorticity field and the corresponding lift coefficient, which are to be estimated by assimilating synthetic surface pressure measurements.